# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ XII Conclusion
In one of the cases tested, $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ turned out to be positive. The analytical approximation then predicts that $R_{,v}$ and $R_{,u}$ should be negative throughout region 3. Therefore $R$ decreases to zero and a spacelike $R=0$ singularity should form. The numerical simulation confirmed this prediction. We presented the spacetime diagram in this case (see Figure XI.15) illustrating the formation of an $R=0$ singularity.
@@ Acknowledgements.
XIII Appendix A - irregularity of “naive $g_{1}$ metric”
@@
$ds^{2}=\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)du\;dv+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (XIII.1)
@@
First of all, the determinant of the metric Eq. (XIII.1) is zero at $r=r_{+}(v)$ because
$1-\frac{2M(v)}{r_{+}(v)}+\frac{Q(v)^{2}}{r_{+}(v)^{2}}=0$ (XIII.2)
@@
$\left.\frac{\partial}{\partial r}\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\right|_{r=r+}=\frac{2M}{r_{+}^{2}}-\frac{2Q^{2}}{r_{+}^{3}}=2\kappa_{+}^{v}(v)$ (XIII.3)
@@
$ds^{2}=2\kappa_{+}^{v}(v)\left(r-r_{+}(v)\right)du\;dv+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (XIII.4)
Around $r_{+},$along a $v=const$ line, $r_{*}$is approximated by (using Eq. (VI.5)):
$\frac{u+v}{2}\equiv r_{*}=\frac{1}{2\kappa_{+}^{v}(v)}\log\left|r-r_{+}\right|+const(v)$ (XIII.5)
@@
$r-r_{+}=e^{2\kappa_{+}^{v}(v)r_{*}}const(v)=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}const(v)$ (XIII.6)
@@
$g_{uv}=\kappa_{+}^{v}(v)\left(r-r_{+}\right)=\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)$ (XIII.8)
@@
$ds^{2}=2\kappa_{+}^{v}(v)h_{0}(v)\;e^{\kappa_{+}^{v}(v)v}dv\;e^{\kappa_{+}^{v}(v)u}du+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (XIII.9)
@@
$ds^{2}=2h_{0}(v)\;\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)u}du\;dV+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (XIII.11)
@@
We’ll now build the Kruskal coordinate $U$. $\frac{\partial U}{\partial u}$ needs to absorb the problematic term $e^{\kappa_{+}^{v}(v)u}$ in the metric. That means that $\frac{\partial U}{\partial u}=C(v)e^{\kappa_{+}^{v}(v)u}$ for some coefficient $C(v)$. From this we get by integration that $U=\frac{C(v)}{\kappa_{+}^{v}(v)}e^{\kappa_{+}^{v}(v)u}+Const(v)$. It turns out that both coefficients $C(v)$ and $Const(v)$ don’t effect the regularity, so we choose
$U\equiv e^{\kappa_{+}(v)u}$ (XIII.12)
@@
$\frac{\partial U}{\partial u}=\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)u}$ (XIII.13)
@@
$\frac{\partial u}{\partial V}=\frac{-1}{\kappa_{+}^{v}(v)^{2}}\;\frac{d}{dV}\left(\kappa_{+}^{v}(v)\right)\log U=h_{1}(V)\log U$ (XIII.15)
@@
$h_{1}(V)\equiv\frac{-1}{\kappa_{+}^{v}(v)^{2}}\;\frac{d}{dV}\left(\kappa_{+}^{v}(v)\right)$ (XIII.16)
@@
$\widehat{g}_{VV}$ $=g_{uV}\frac{\partial u}{\partial V}\frac{\partial V}{\partial V}=\kappa_{+}^{v}(v)h_{0}(v)Uh_{1}(V)\log U$ (XIII.17)
$=\kappa_{+}^{v}(v)h_{0}(v)h_{1}(V)U\log U=h_{2}(v)U\log U$ (XIII.18)
@@
$\widehat{g}_{VV,U}=h_{2}(v)\left(\log U+1\right)$ (XIII.20)
@@
$\widehat{g}_{UV}=g_{uV}\frac{\partial u}{\partial U}\frac{\partial V}{\partial V}=\kappa_{+}^{v}(v)h_{0}(v)U\frac{1}{U\kappa_{+}^{v}(v)}=h_{0}(v)$ (XIII.21)
@@
$ds^{2}=h_{2}(V)U\log UdV^{2}+2h_{0}(v)dU\;dV+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (XIII.22)
@@
$R_{curvature}$ $=\frac{h_{2}(v)}{Uh_{0}(v)^{2}}+\frac{4h_{2}(v)r\;r_{,U}\log U}{r^{2}h_{0}(v)^{2}}$ (XIII.23)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2h_{0}(v)^{2}+4h_{2}(v)r\;r_{,U}-4h_{0}(v)r_{,V}r_{,U}+\right)$ (XIII.24)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2Uh_{2}(v)\kappa_{+}(v)r_{,U}^{2}\log U-8h_{0}(v)r\;r_{,UV}+4Uh_{2}(v)r\;r_{,UU}\log U\right)$ (XIII.25)
We wish to show that this expression diverges as $U\rightarrow 0$ at the EH. To do that, we calculate from Eq. (XIII.7) (the approximation of $r-r_{+}$ near the EH) the derivatives of $r$:
$r-r_{+}=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)\equiv Uh_{3}(v)$ (XIII.26)
@@
$R_{curvature}$ $=\frac{h_{2}(v)}{Uh_{0}(v)^{2}}+\frac{4h_{2}(v)rh_{3}(v)\log U}{r^{2}h_{0}(v)^{2}}$ (XIII.31)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2h_{0}(v)^{2}+4h_{2}(v)rh_{3}(v)-4h_{0}(v)Uh_{3}^{\prime}(v)h_{3}(v)\right)$ (XIII.32)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2Uh_{2}(v)\kappa_{+}(v)h_{3}(v)^{2}\log U-8h_{0}(v)rh_{3}^{\prime}(v)\right)$ (XIII.33)
@@
XIV Appendix B - determination of the initial values from $R_{,u}$, $S_{,u}$
We know the geometry in the near EH region 1, it is given by the $g_{1}$ metric Eq. (VI.29,VI.30). We use this metric to construct the initial conditions for the numerically evolving metric in region 2 (and subsequently region 3). In principle, this requires the initial values of the unknowns $R$,$S$ and their time derivatives on a hypersurface of constant $T\equiv\frac{u+v}{2}$, namely on $T=T_{0}$. We denote $Y\equiv\frac{v-u}{2}$. Thus the standard initial value problem would in principle require the four functions $R$, $R_{,T}$, $S$, $S_{,T}$ to produce a unique solution (where the partial derivatives are taken while keeping $Y$ fixed). Because $R_{,u}=\frac{1}{2}R_{,T}-\frac{1}{2}R_{,Y}$, and $R_{,Y}$ is in principle known along the initial hypersurface (from the initial function $R(Y)$), we can also use $R_{,u}$ for the initial values instead of $R_{,T}$. We prefer using $R_{,u}$ instead of $R_{,T}$ because it’s easier to calculate, as $M(v)$ is constant along lines of constant $v.$
In practice, our numerical algorithm requires two adjacent rows of initial values of $R$ and $S$ (but not of $R_{,T}$ and $S_{,T}$), namely the row $T=T_{0}$ and the subsequent row. We show in Figure XIV.1 the two rows of initial values, $T=T_{0}$ in blue and the next one in red, which are required for evolution into the black points. We are given $R_{,u}$ and $S_{,u}$ at $T=T_{0}$ directly from the initial values of $g_{1}$ metric, and need to translate it into values of $R$ and $S$ on the red row.
Figure XIV.1: Initial values rows
Figure XIV.2: Initial values points
We proceed to calculate $R$ at the red point $(i,j)$ as shown in Figure XIV.2 by calculating $R_{,u}$ at the green point $(i,j-\nicefrac{{1}}{{2}})$, and use that, and the value of $R$ at the blue point (i,j1)(i,j-1), to obtain $R$ at the red grid point $(i,j)$ (see Eq. (XIV.3) below). To obtain $R_{,u}(i,j-\nicefrac{{1}}{{2}})$, we first calculate $R_{,u}$ at the yellow point $R(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ by averaging:
$R_{,u}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})=\frac{1}{2}\left(R_{,u}(i-1,j)+R_{,u}(i,j-1)\right)$ (XIV.1)
@@ XIV Appendix B - determination of the initial values from $R_{,u}$, $S_{,u}$
$R_{,u}(i,j-\nicefrac{{1}}{{2}})=R_{,u}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})+\frac{1}{2}dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (XIV.2)
@@
$R(i,j)=R(i,j-1)+du\left(\frac{1}{2}\left(R_{,u}(i-1,j)+R_{,u}(i,j-1)\right)+\frac{1}{2}dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})\right)$ (XIV.4)

### v2

@@ 12 Conclusion
In one of the cases tested, $\widetilde{T}_{vv}^{(-)}$ and $\widetilde{T}_{uu}^{(-)}$ turned out to be positive. The analytical approximation then predicts that $R_{,v}$ and $R_{,u}$ should be negative throughout region 3. Therefore $R$ decreases to zero and a spacelike $R=0$ singularity should form. The numerical simulation confirmed this prediction. We presented the spacetime diagram in this case (see Figure 11.15) illustrating the formation of an $R=0$ singularity.
@@ 12.1 Acknoledgments
13 Appendix A - irregularity of “naive $g_{1}$ metric”
@@
$ds^{2}=\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)du\;dv+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.1)
@@
First of all, the determinant of the metric Eq. (13.1) is zero at $r=r_{+}(v)$ because
$1-\frac{2M(v)}{r_{+}(v)}+\frac{Q(v)^{2}}{r_{+}(v)^{2}}=0$ (13.2)
@@
$\left.\frac{\partial}{\partial r}\left(1-\frac{2M(v)}{r}+\frac{Q(v)^{2}}{r^{2}}\right)\right|_{r=r+}=\frac{2M}{r_{+}^{2}}-\frac{2Q^{2}}{r_{+}^{3}}=2\kappa_{+}^{v}(v)$ (13.3)
@@
$ds^{2}=2\kappa_{+}^{v}(v)\left(r-r_{+}(v)\right)du\;dv+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.4)
Around $r_{+},$along a $v=const$ line, $r_{*}$is approximated by (using Eq. (6.5)):
$\frac{u+v}{2}\equiv r_{*}=\frac{1}{2\kappa_{+}^{v}(v)}\log\left|r-r_{+}\right|+const(v)$ (13.5)
@@
$r-r_{+}=e^{2\kappa_{+}^{v}(v)r_{*}}const(v)=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}const(v)$ (13.6)
@@
$g_{uv}=\kappa_{+}^{v}(v)\left(r-r_{+}\right)=\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)$ (13.8)
@@
$ds^{2}=2\kappa_{+}^{v}(v)h_{0}(v)\;e^{\kappa_{+}^{v}(v)v}dv\;e^{\kappa_{+}^{v}(v)u}du+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.9)
@@
$ds^{2}=2h_{0}(v)\;\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)u}du\;dV+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.11)
@@
We’ll now build the Kruskal coordinate $U$. $\frac{\partial U}{\partial u}$ needs to absorb the problematic term $e^{\kappa_{+}^{v}(v)u}$ in the metric. That means that $\frac{\partial U}{\partial u}=C(v)e^{\kappa_{+}^{v}(v)u}$ for some coefficient $C(v)$. From this we get by integration that $U=\frac{C(v)}{\kappa_{+}^{v}(v)}e^{\kappa_{+}^{v}(v)u}+Const(v)$. It turns out that both coefficients $C(v)$ and $Const(v)$ do not affect the regularity, so we choose
$U\equiv e^{\kappa_{+}(v)u}$ (13.12)
@@
$\frac{\partial U}{\partial u}=\kappa_{+}^{v}(v)e^{\kappa_{+}^{v}(v)u}$ (13.13)
@@
$\frac{\partial u}{\partial V}=\frac{-1}{\kappa_{+}^{v}(v)^{2}}\;\frac{d}{dV}\left(\kappa_{+}^{v}(v)\right)\log U=h_{1}(V)\log U$ (13.15)
@@
$h_{1}(V)\equiv\frac{-1}{\kappa_{+}^{v}(v)^{2}}\;\frac{d}{dV}\left(\kappa_{+}^{v}(v)\right)$ (13.16)
@@
$\widehat{g}_{VV}$ $=g_{uV}\frac{\partial u}{\partial V}\frac{\partial V}{\partial V}=\kappa_{+}^{v}(v)h_{0}(v)Uh_{1}(V)\log U$ (13.17)
$=\kappa_{+}^{v}(v)h_{0}(v)h_{1}(V)U\log U=h_{2}(v)U\log U$ (13.18)
@@
$\widehat{g}_{VV,U}=h_{2}(v)\left(\log U+1\right)$ (13.20)
@@
$\widehat{g}_{UV}=g_{uV}\frac{\partial u}{\partial U}\frac{\partial V}{\partial V}=\kappa_{+}^{v}(v)h_{0}(v)U\frac{1}{U\kappa_{+}^{v}(v)}=h_{0}(v)$ (13.21)
@@
$ds^{2}=h_{2}(V)U\log UdV^{2}+2h_{0}(v)dU\;dV+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)$ (13.22)
@@
$R_{curvature}$ $=\frac{h_{2}(v)}{Uh_{0}(v)^{2}}+\frac{4h_{2}(v)r\;r_{,U}\log U}{r^{2}h_{0}(v)^{2}}$ (13.23)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2h_{0}(v)^{2}+4h_{2}(v)r\;r_{,U}-4h_{0}(v)r_{,V}r_{,U}+\right)$ (13.24)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2Uh_{2}(v)\kappa_{+}(v)r_{,U}^{2}\log U-8h_{0}(v)r\;r_{,UV}+4Uh_{2}(v)r\;r_{,UU}\log U\right)$ (13.25)
We wish to show that this expression diverges as $U\rightarrow 0$ at the EH. To do that, we calculate from Eq. (13.7) (the approximation of $r-r_{+}$ near the EH) the derivatives of $r$:
$r-r_{+}=e^{\kappa_{+}^{v}(v)v+\kappa_{+}^{v}(v)u}h_{0}(v)\equiv Uh_{3}(v)$ (13.26)
@@
$R_{curvature}$ $=\frac{h_{2}(v)}{Uh_{0}(v)^{2}}+\frac{4h_{2}(v)rh_{3}(v)\log U}{r^{2}h_{0}(v)^{2}}$ (13.31)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2h_{0}(v)^{2}+4h_{2}(v)rh_{3}(v)-4h_{0}(v)Uh_{3}^{\prime}(v)h_{3}(v)\right)$ (13.32)
$+\frac{1}{r^{2}h_{0}(v)^{2}}\left(2Uh_{2}(v)\kappa_{+}(v)h_{3}(v)^{2}\log U-8h_{0}(v)rh_{3}^{\prime}(v)\right)$ (13.33)
@@
14 Appendix B - determination of the initial values from $R_{,u}$, $S_{,u}$
We know the geometry in the near EH region 1, it is given by the $g_{1}$ metric Eq. (6.29,6.30). We use this metric to construct the initial conditions for the numerically evolving metric in region 2 (and subsequently region 3). In principle, this requires the initial values of the unknowns $R$,$S$ and their time derivatives on a hypersurface of constant $T\equiv\frac{u+v}{2}$, namely on $T=T_{0}$. We denote $Y\equiv\frac{v-u}{2}$. Thus the standard initial value problem would in principle require the four functions $R$, $R_{,T}$, $S$, $S_{,T}$ to produce a unique solution (where the partial derivatives are taken while keeping $Y$ fixed). Because $R_{,u}=\frac{1}{2}R_{,T}-\frac{1}{2}R_{,Y}$, and $R_{,Y}$ is in principle known along the initial hypersurface (from the initial function $R(Y)$), we can also use $R_{,u}$ for the initial values instead of $R_{,T}$. We prefer using $R_{,u}$ instead of $R_{,T}$ because it’s easier to calculate, as $M(v)$ is constant along lines of constant $v.$
In practice, our numerical algorithm requires two adjacent rows of initial values of $R$ and $S$ (but not of $R_{,T}$ and $S_{,T}$), namely the row $T=T_{0}$ and the subsequent row. We show in Figure 14.1 the two rows of initial values, $T=T_{0}$ in blue and the next one in red, which are required for evolution into the black points. We are given $R_{,u}$ and $S_{,u}$ at $T=T_{0}$ directly from the initial values of $g_{1}$ metric, and need to translate it into values of $R$ and $S$ on the red row.
Figure 14.1: Initial values rows
Figure 14.2: Initial values points
We proceed to calculate $R$ at the red point $(i,j)$ as shown in Figure 14.2 by calculating $R_{,u}$ at the green point $(i,j-\nicefrac{{1}}{{2}})$, and use that, and the value of $R$ at the blue point (i,j1)(i,j-1), to obtain $R$ at the red grid point $(i,j)$ (see Eq. (14.3) below). To obtain $R_{,u}(i,j-\nicefrac{{1}}{{2}})$, we first calculate $R_{,u}$ at the yellow point $R(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ by averaging:
$R_{,u}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})=\frac{1}{2}\left(R_{,u}(i-1,j)+R_{,u}(i,j-1)\right)$ (14.1)
@@ 14 Appendix B - determination of the initial values from $R_{,u}$, $S_{,u}$
$R_{,u}(i,j-\nicefrac{{1}}{{2}})=R_{,u}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})+\frac{1}{2}dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})$ (14.2)
@@
$R(i,j)=R(i,j-1)+du\left(\frac{1}{2}\left(R_{,u}(i-1,j)+R_{,u}(i,j-1)\right)+\frac{1}{2}dvR_{,uv}(i-\nicefrac{{1}}{{2}},j-\nicefrac{{1}}{{2}})\right)$ (14.4)